# Ion Temperature Anisotropy Limits from Magnetic Curvature Scattering in Magnetotail Reconnection Jets

### v1

@@ III.2 Lunar distances
To complement MMS observations, we use data from the Acceleration, Reconnection, Turbulence, and Electrodynamics of the Moon’s Interaction with the Sun (ARTEMIS) mission. ARTEMIS consists of two spacecraft orbiting the Moon, enabling measurements of the magnetotail at lunar distances ($X_{GSM}\sim-60R_{E}$). The magnetotail at such distances often exhibits mixed plasma compositions, with both hot and suprathermal ($\gtrsim 1$ keV) ions originating from near-Earth and distant-tail reconnection, as well as cold (<100<100 eV) ions of ionospheric or mantle origin [wang_properties_2014, runov_properties_2023]. At these distances, the dipole contribution to the equatorial magnetic field is essentially nonexistent, allowing for a significant fraction of demagnetized (hot) ions executing Speiser-like motion to coexist with magnetized cold ions undergoing $\mathbf{E}\times\mathbf{B}$ motion [kamaletdinov_ion_2025]. Thus, the ARTEMIS dataset offers a unique opportunity to study both the effects of Speiser-like ions and the Hall electric field that governs cold-ion dynamics. The resultant dataset includes 1261 magnetotail current sheet crossings collected between 2012 and 2022 [kamaletdinov_thin_2024], spanning $X_{GSM}\in[-65,-50]R_{E}$. Magnetic field measurements are provided by the fluxgate magnetometer [auster_themis_2008], and ion parallel and perpendicular temperatures are measured by electrostatic analyzers [mcfadden_themis_2008].
We plot the joint PDF of $\beta_{i\parallel}$ and $T_{i\bot}/T_{i\parallel}$ in Fig. 4(b). As expected, we find that the data mostly fall at higher-$\beta_{i\parallel}$ compared with near-Earth magnetotail [Fig. 4(a)], due to the weaker magnetic field and larger temperature, and nearly isotropic ion VDFs $T_{i\bot}/T_{i\parallel}\approx 1$. Although the data are very concentrated in the high $\beta_{i\parallel}$ and $T_{i\bot}/T_{i\parallel}\approx 1$ end, the distribution resembles that of MMS data [Fig. 4a], and the iso contours also qualitatively resemble the thresholds, providing further support for these limits.
@@ IV Numerical simulations
To further validate the limits for the ion anisotropy obtained in Eqs. 8 and 12, we analyze two fully kinetic particle-in-cell (PIC) simulations of symmetric magnetic reconnection in a magnetotail-like configuration [an_configuration_2022, norgren_presence_2021].
@@ IV.1 Lembège–Pellat equilibrium
The first simulation from Ref. an_configuration_2022, hereafter PIC-LP, employs a generalized Lembège–Pellat current sheet equilibrium [lembege_stability_1982, lu_hall_2019], which includes a background plasma population and a polarization electric field, typical of magnetotail current sheets [artemyev_thin_2009, runov_local_2006].
@@ IV.1.1 Numerical setup
We perform a PIC simulation [pritchett_kinetic_2001, pritchett_externally_2005] of driven magnetic reconnection starting from the Lembège–Pellat equilibrium described below in Section IV.1.2. The simulation is two-dimensional in configuration space (2D) and three-dimensional in velocity space (3V). Normalizations are based on ion quantities: time is normalized to the inverse ion cyclotron frequency $\omega_{ci}=eB_{0}/m_{i}$, lengths to the ion inertial length $d_{i}=\sqrt{m_{i}/\mu_{0}n_{0}e^{2}}$, velocities to the Alfvén speed $V_{A0}$, and energies to $m_{i}V_{A0}^{2}$. The simulation domain extends over $L_{z}/2,L_{z}/2]\times[-L_{x},0]$, with $L_{x}=64\,d_{i}$ and $L_{z}=16\,d_{i}$, and a grid cell size of $\delta/d_{i}=1/32$. The Debye length is $\lambda_{De}=0.46\,\delta$, the ion-to-electron mass ratio is $m_{i}/m_{e}=100$, and the reduced speed of light is $c/V_{A0}=20$. The time step is $\Delta t=0.001\,\omega_{ci}^{-1}$. The reference density $n_{0}$ is represented by 425 particles per cell, corresponding to a total of $\sim 2\times 10^{8}$ particles initially. The simulation domain has open boundary conditions at $x/d_{i}=-L_{x}$ and $x/d_{i}=0$ that allow particles to exit the domain while continuously injecting back particles in the system [aldrich_particle_1985, pritchett_externally_2005], and reflective boundary conditions at $z/d_{i}=-L_{z}/2$ and $z/d_{i}=L_{z}/2$.
@@ IV.1.2 Initial equilibrium
$\dfrac{\partial^{2}A_{y}}{\partial z^{2}}=-\mu_{0}\sum_{\alpha}q_{\alpha}n_{0}v_{\alpha D}\exp\left(-\dfrac{q_{\alpha}\phi}{T_{\alpha 0}}+\dfrac{q_{\alpha}v_{\alpha D}A_{y}}{T_{\alpha 0}}\right),$ (13)
$\sum_{\alpha}q_{\alpha}n_{0}\exp\left(-\dfrac{q_{\alpha}\phi}{T_{\alpha 0}}+\dfrac{q_{\alpha}v_{\alpha D}A_{y}}{T_{\alpha 0}}\right)\\ +\sum_{\alpha}q_{\alpha}n_{b}\exp\left(-\dfrac{q_{\alpha}\phi}{T_{\alpha b}}\right)=0,$ (14)
where we have separated the background population from the current-carrying population in the current sheet. Here, $\alpha=i,e$ denotes the particle species, $q_{\alpha}$ the particle charge, $n_{0}$ and $n_{b}$ the number densities of the current sheet and background populations, respectively, $T_{\alpha 0}$ and $T_{\alpha b}$ the current sheet and background species temperatures, and $v_{\alpha D}$ the drift speed. We use $n_{b}/n_{0}=1/5$, $T_{i0}=T_{ib}=5m_{i}V_{A0}^{2}/12$, $T_{e0}=T_{eb}=m_{i}V_{A0}^{2}/12=T_{i0}/5$, $v_{iD}=V_{A0}/3$ and $v_{eD}=-5V_{A0}/3$. Solving Eqs. 13 and 14, with the number densities and current densities calculated using a drifting Maxwellian equilibrium VDF [lembege_stability_1982], and the boundary conditions at the current sheet center $z=0$ where $B_{x}=0$ and $B_{z}=\varepsilon B_{0}$:
@@
This initial configuration, featuring a finite magnetic field component normal to the current sheet ($B_{z}\neq 0$), is stable against spontaneous reconnection [pellat_does_1991]. We initiate reconnection by imposing electric fields:
$E_{y,drive}(x,z=\pm L_{z}/2,t)=E_{y0}f(t)\sin^{2}\left(\frac{\pi x}{L_{x}}\right)$ (16)
at the $z/d_{i}=\pm L_{z}/2$ boundaries. Here, $E_{y0}/E_{0}=0.4$, and the $f(t)=\tanh\left(\omega t\right)\operatorname{sech}^{2}\left(\omega t\right)$ with the characteristic frequency $\omega/\omega_{ci}=20$. This forcing drives plasma inflows that compress and thin the current sheet, therefore forcing magnetic reconnection.
@@ IV.1.3 Results
Figure 3(i) shows the ion anisotropy $T_{i\bot}/T_{i\parallel}$ at $t\omega_{ci}=110$, slightly after the peak of the reconnection rate, which occurs at $t\omega_{ci}=84$ (see Figure 1 in Ref. an_configuration_2022). Here, we calculate $T_{i\bot}$ and $T_{i\parallel}$ using the rotation to the MFA as described in Section III. We note that the extreme parallel anisotropies at the top and bottom of the domain arise from the reflective boundary conditions and are not physical. In the reconnection outflows, the parallel temperature dominates outside the current sheet, whereas the perpendicular temperature dominates at the current sheet center. In particular, we compute the adiabaticity parameter $\kappa$ [Fig. 3(j)], and overlay the contour of $\kappa=1$ in Fig. 3(i). We find that the contour of $\kappa=1$ clearly separates the parallel anisotropy outside the current sheet, where $\kappa>1$, and the perpendicular anisotropy inside the current sheet, where $\kappa\lesssim 1$. This is consistent with the MMS observations [Fig. 2], and further supports our model.
@@
We present the joint PDF of $\beta_{i\parallel}$ and $T_{i\bot}/T_{i\parallel})$ in the PIC-LP simulation in Fig. 4(c). To compare the PIC-LP simulation with MMS observations, we restrict to regions in the outflow where $\beta_{i}>0.1$ and $V_{ix}>0$ (green contour in Figs. 3(i) and 3(j)). To collect a statistically significant number of points, we combine data within $t\omega\in[84,128]$ during which the reconnection electric field $E_{y,rec}$ at the X-line monotonically decreases from its peak value $E_{y,rec}^{max}/E_{0}=0.07$ at $t\omega_{ci}=84$ but remains $E_{y,rec}/E_{y,rec}^{max}>1/2$. We find that the distribution of ion anisotropy in the PIC-LP simulation is relatively similar to that observed by MMS. However, we find only weak parallel anisotropy, possibly due to the initially hot background with $T_{ib}=T_{i0}$ and the absence of initially cold ions. Nevertheless, we find that the threshold $\mathcal{R}_{s}$ bounds the distribution well, supporting the derived thresholds.
@@ IV.2 Harris equilibrium with cold ions
The second simulation from Ref. norgren_presence_2021, hereafter referred to as PIC-HC, employs a Harris current sheet equilibrium with cold ions as the inflow populations. While this configuration is slightly more idealized and suited to study anti-parallel magnetic reconnection, the inclusion of initially cold ions provides the opportunity to test the limit for parallel anisotropy.
@@ IV.2.1 Numerical setup
We perform a particle-in-cell (PIC) simulation [hesse_diffusion_1999] of driven magnetic reconnection starting from the Harris equilibrium described in Section IV.2.2. The simulation is 2D in configuration space and 3V in velocity space. Normalizations are those of the simulation of the PIC-LP run. The simulation domain extends over $L_{z}/2,L_{z}/2]\times[0,L_{x}]$, with $L_{x}=200\,d_{i}$ and $L_{z}=12.5\,d_{i}$, divided into a grid of $6400\times 1600$ cells. The ion-to-electron mass ratio is $m_{i}/m_{e}=100$, and the reduced speed of light is $c/V_{A0}=20$. The time step is $\Delta t=0.5\omega_{pe}^{-1}=0.0025\omega_{ci}^{-1}$. The inflow density $n_{c}$ is represented by $30\sim$ particles per cell, and the total number of particles is $\sim 4.8\times 10^{8}$. The simulation domain has periodic boundary conditions at $x/d_{i}=0$ and $x/d_{i}=L_{x}$, and reflective boundary conditions at $z/d_{i}=-L_{z}/2$ and $z/d_{i}=L_{z}/2$.
@@ IV.2.2 Initial equilibrium
$n=n_{0}\operatorname{sech}^{2}\left(\frac{z}{L}\right)+n_{c}\left[\frac{1}{2}+\frac{1}{2}\tanh\left(\frac{|z|-2L}{L/2}\right)\right],$ (18)
@@ IV.2.3 Results
Figure 5: Anisotropy $T_{i\bot}/T_{i\parallel}$ (a) and adiabaticity parameter $\kappa$ (b) of initially cold ions in the PIC-HC simulation. The black contour indicates $\kappa=1$. The green contour indicates the region of interest for Fig. 4(d).
We present the anisotropy $T_{i\bot}/T_{i\parallel}$ of the initially cold ions at time $t=120~\omega_{ci}^{-1}$ in Fig. 5. At the time and place of investigation, the effect of the periodicity of the simulation is negligible, and the reconnection exhaust is completely dominated by the initially cold ions [norgren_presence_2021]. Similar to that seen in Fig. 3(j), the temperature anisotropy is mainly parallel outside the current sheet, except in the inflow near the X-line, near the dipolarization front, and at ion mirror points. We compute the adiabaticity parameter $\kappa$ [Fig. 5(b)], and overlay the contour of $\kappa=1$ in Fig. 5(a). We find that, similar to that in the PIC-LP run [Fig. 3(i)], the anisotropy tends to be larger in regions where $\kappa\lesssim 1$ than in the vicinity outside of the current sheet where $\kappa>1$.
We present the joint PDF of $\beta_{i\parallel}$ and $T_{i\bot}/T_{i\parallel}$ in Fig. 4(d). Similar to the PIC-LP run and MMS observations, we restrict to regions of the outflow where $\beta_{i}>0.1$ and $V_{ix}>0$ (green contour in Fig. 5). We find that, similar to MMS observations, the thresholds limit the distribution relatively well for $T_{i\bot}/T_{i\parallel}<1$. In contrast, the perpendicular anisotropy slightly exceeds the marginal stability condition $\kappa=1$. However, similar to MMS observations, this may be attributed to overestimates of $\eta_{e}$ and $\eta_{s}$, or weak scattering, i.e., slow growth of the instability.
@@ V Discussion
We derived two limits for the ion anisotropy in a magnetotail-like quasi-1D current sheet with a finite normal magnetic field, incorporating three ion populations: an adiabatic cold population of two counter-streaming beams, an isotropic hot population, and a quasi-adiabatic Speiser population. Here, we discuss the physical interpretation of these limits and their implications for the stability of the current sheet.
@@
The first limit [Eq. 8], which concerns the parallel anisotropy, resembles the firehose condition, in which the plasma pressure force balances the magnetic tension. In the particular case where $\eta_{c}=1$, i.e., $T_{c\bot}=T_{i\parallel}$ so that the ion VDF consists of a single population with anisotropy $T_{i\bot}/T_{i\parallel}$, the threshold reduces to the well-known fluid firehose condition [cowley_plasma_1980]. In contrast, in the limit $\eta_{c}^{2}\beta_{i\parallel}\rightarrow 0$, the threshold becomes $\mathcal{R}_{c}(\eta_{c}\rightarrow 0,\beta_{i\parallel})\rightarrow\eta_{c}^{2}\beta_{i\parallel}$. This regime corresponds to strongly magnetized cold ions, for which $T_{c\bot}/B^{2}\ll 1$, implying that larger magnetic field curvature is required to scatter the cold ions. Consequently, cold ions can maintain force balance even in highly stretched magnetic field configurations.
@@
The second limit [Eq. 12], which concerns the perpendicular anisotropy, defines the threshold above which the hot isotropic ion population is scattered into quasi-adiabatic transient Speiser orbits. Since the adiabaticity parameter for one particle inversely scales with the particle’s perpendicular velocity $\kappa\propto 1/\sqrt{v_{\bot}}$, for a given increase of the magnetic curvature due to e.g., current sheet thinning/stretching, the supra-thermal ions with the corresponding perpendicular velocity, initially in the hot isotropic background, will be scattered into Speiser orbits. As a result, the number density $n_{h}$ and temperature $T_{h}$ of the hot isotropic background will decrease, and concurrently, the number density $n_{s}$ of the quasi-adiabatic Speiser population will increase, thereby increasing their current density. Eventually, the resulting ion VDF will be composed of a core plus a dense drifting beam Speiser population. If the drift velocity is sufficiently large compared with the thermal speed and the drifting beam is sufficiently dense with respect to the core, this will lead to the growth of a drift-kink type instability [lapenta_kinetic_1997, daughton_unstable_1999, zelenyi_low_2009].
@@
The drift-kink instability may act back on the ion VDFs, scattering quasi-adiabatic Speiser ions into the hot isotropic population. Recent in situ observations suggest that the growth of the drift-kink instability may ultimately lead to a thickening of the current sheet [richard_observations_2021]. During the early stage of the drift-kink instability, the enhancement of the quasi-adiabatic Speiser ion current causes the current sheet to thin, generating strong density gradients. These density gradients result in the growth of the lower-hybrid drift instability (LHDI) [daughton_nonlinear_2002, daughton_electromagnetic_2003, daughton_nonlinear_2004], which is localized at the current sheet edges due to the strong stabilization effect of finite $\beta$ [davidson_effects_1977]. As demonstrated by Ref. daughton_nonlinear_2002 and daughton_nonlinear_2004, the nonlinear evolution of the LHDI induces resonant scattering of crossing Speiser ions into the non-crossing region of phase space. This process leads to a net negative charge within the current sheet and the formation of an electrostatic potential across it, which in turn results in a thickening of the current sheet [sitnov_structure_2006]. As the radius of curvature increases due to the thickening of the current sheet, Speiser ions are subsequently scattered back into the hot, isotropic ion population, thereby relaxing the ion VDFs toward equilibrium.
@@
We show that, to maintain the stability of the current sheet, the ion anisotropy is constrained by two limits set by curvature scattering. Nevertheless, as recent theoretical developments and in situ observations suggest, plasma anisotropy plays a crucial role in the force balance of magnetotail-like current sheets [artemyev_configuration_2021, an_configuration_2022, egedal_plasma_2023]. Since electrons are generally isotropic, their pressure force accounts for only up to 10%-30% of the magnetic tension [artemyev_contribution_2019]. Recent numerical simulations have further revealed that even a small population of Speiser ions can contribute significantly to maintaining force balance through their pressure contribution [arnold_pic_2023]. Therefore, our results indicate that Speiser ions play a crucial role in maintaining the force balance of the current sheet and consistently determine the magnetic field curvature, i.e., the degree of stretching of the current sheet, necessary to ensure its stability.

### v2

@@ III.2 Lunar distances
To complement MMS observations, we use data from the Acceleration, Reconnection, Turbulence, and Electrodynamics of the Moon’s Interaction with the Sun (ARTEMIS) mission. ARTEMIS consists of two spacecraft orbiting the Moon, enabling measurements of the magnetotail at lunar distances ($X_{GSM}\sim-60R_{E}$). The magnetotail at such distances often exhibits mixed plasma compositions, with both hot and suprathermal ($\gtrsim 1$ keV) ions originating from near-Earth and distant-tail reconnection, as well as cold (<100<100 eV) ions of ionospheric or mantle origin [73, 66]. At these distances, the dipole contribution to the equatorial magnetic field is essentially nonexistent, allowing for a significant fraction of demagnetized (hot) ions executing Speiser-like motion to coexist with magnetized cold ions undergoing $\mathbf{E}\times\mathbf{B}$ motion [43]. Thus, the ARTEMIS dataset offers a unique opportunity to study both the effects of Speiser-like ions and the Hall electric field that governs cold-ion dynamics. The resultant dataset includes 1261 magnetotail current sheet crossings collected between 2012 and 2022 [42], spanning $X_{GSM}\in[-65,-50]R_{E}$. Magnetic field measurements are provided by the fluxgate magnetometer [14], and ion parallel and perpendicular temperatures are measured by electrostatic analyzers [52].
We plot the joint PDF of $\beta_{i\parallel}$ and $T_{i\bot}/T_{i\parallel}$ in Fig. 4(b). We find that the data mostly fall at higher-$\beta_{i\parallel}$ compared with near-Earth magnetotail [Fig. 4(a)], due to the weaker magnetic field and larger temperature, and nearly isotropic ion VDFs $T_{i\bot}/T_{i\parallel}\approx 1$. Although the data are very concentrated in the high $\beta_{i\parallel}$ and $T_{i\bot}/T_{i\parallel}\approx 1$ end, the distribution resembles that of MMS data [Fig. 4a], and the iso-contours also qualitatively resemble the thresholds, providing further support for these limits.
@@ IV Numerical simulations
To further validate the limits for the ion anisotropy obtained in Eqs. 7 and 11, we analyze two fully kinetic particle-in-cell (PIC) simulations of symmetric magnetic reconnection in a magnetotail-like configuration [3, 55].
@@ IV.1 Lembège–Pellat equilibrium
The first simulation from Ref. 3, hereafter PIC-LP, employs a generalized Lembège–Pellat current sheet equilibrium [46, 50], which includes a background plasma population and a polarization electric field, typical of magnetotail current sheets [11, 67].
@@ IV.1.1 Numerical setup
We perform a PIC simulation [59, 60] of driven magnetic reconnection starting from the Lembège–Pellat equilibrium described below in Section IV.1.2. The simulation is two-dimensional in configuration space (2D) and three-dimensional in velocity space (3V). Normalizations are based on ion quantities: time is normalized to the inverse ion cyclotron frequency $\omega_{ci}=eB_{0}/m_{i}$, lengths to the ion inertial length $d_{i}=\sqrt{m_{i}/\mu_{0}n_{0}e^{2}}$, velocities to the Alfvén speed $V_{A0}$, and energies to $m_{i}V_{A0}^{2}$. The simulation domain extends over $L_{z}/2,L_{z}/2]\times[-L_{x},0]$, with $L_{x}=64\,d_{i}$ and $L_{z}=16\,d_{i}$, and a grid cell size of $\delta/d_{i}=1/32$. The Debye length is $\lambda_{De}=0.46\,\delta$, the ion-to-electron mass ratio is $m_{i}/m_{e}=100$, and the reduced speed of light is $c/V_{A0}=20$. The time step is $\Delta t=0.001\,\omega_{ci}^{-1}$. The reference density $n_{0}$ is represented by 425 particles per cell, corresponding to a total of $\sim 2\times 10^{8}$ particles initially. The simulation domain has open boundary conditions at $x/d_{i}=-L_{x}$ and $x/d_{i}=0$ that allow particles to exit the domain while continuously injecting back particles in the system [1, 60], and reflective boundary conditions at $z/d_{i}=-L_{z}/2$ and $z/d_{i}=L_{z}/2$.
@@ IV.1.2 Initial equilibrium
[…]
{multline}
[…]
@@
This initial configuration, featuring a finite magnetic field component normal to the current sheet ($B_{z}\neq 0$), is stable against spontaneous reconnection [57]. We initiate reconnection by imposing electric fields:
$E_{y,drive}(x,z=\pm L_{z}/2,t)=E_{y0}f(t)\sin^{2}\left(\frac{\pi x}{L_{x}}\right),$ (14)
[…]
@@ IV.1.3 Results
Figure 3(i) shows the ion anisotropy $T_{i\bot}/T_{i\parallel}$ at $t\omega_{ci}=110$, slightly after the peak of the reconnection rate, which occurs at $t\omega_{ci}=84$ (see Figure 1 in Ref. 3). Here, we calculate $T_{i\bot}$ and $T_{i\parallel}$ using the rotation to the MFA as described in Section III. We note that the extreme parallel anisotropies at the top and bottom of the domain arise from the reflective boundary conditions and are not physical. In the reconnection outflows, the parallel temperature dominates outside the current sheet, whereas the perpendicular temperature dominates at the current sheet center. In particular, we compute the adiabaticity parameter $\kappa$ [Fig. 3(j)], and overlay the contour of $\kappa=1$ in Fig. 3(i). We find that the contour of $\kappa=1$ clearly separates the parallel anisotropy outside the current sheet, where $\kappa>1$, and the perpendicular anisotropy inside the current sheet, where $\kappa\lesssim 1$. This is consistent with the MMS observations [Fig. 2], and further supports our model.
@@
We present the joint PDF of $\beta_{i\parallel}$ and $T_{i\bot}/T_{i\parallel}$ in the PIC-LP simulation in Fig. 4(c). To compare the PIC-LP simulation with MMS observations, we restrict to regions in the outflow where $\beta_{i}>0.1$ and $V_{ix}>0$ (green contour in Figs. 3(i) and 3(j)). To collect a statistically significant number of points, we combine data within $t\omega\in[84,128]$ during which the reconnection electric field $E_{y,rec}$ at the X-line monotonically decreases from its peak value $E_{y,rec}^{max}/E_{0}=0.07$ at $t\omega_{ci}=84$ but remains $E_{y,rec}/E_{y,rec}^{max}>1/2$. We find that the distribution of ion anisotropy in the PIC-LP simulation is relatively similar to that observed by MMS. However, we find only weak parallel anisotropy, possibly due to the initially hot background with $T_{ib}=T_{i0}$ and the absence of initially cold ions. Nevertheless, we find that the threshold $\mathcal{R}_{s}$ bounds the distribution well, supporting the derived thresholds.
@@ IV.2 Harris equilibrium with cold ions
The second simulation from Ref. 55, hereafter referred to as PIC-HC, employs a Harris current sheet equilibrium with cold ions as the inflow populations. While this configuration is slightly more idealized and suited to study anti-parallel magnetic reconnection, the inclusion of initially cold ions provides the opportunity to test the limit for parallel anisotropy.
@@ IV.2.1 Numerical setup
We perform a particle-in-cell (PIC) simulation [38] of driven magnetic reconnection starting from the Harris equilibrium described in Section IV.2.2. The simulation is 2D in configuration space and 3V in velocity space. Normalizations are those of the simulation of the PIC-LP run. The simulation domain extends over $L_{z}/2,L_{z}/2]\times[0,L_{x}]$, with $L_{x}=200\,d_{i}$ and $L_{z}=12.5\,d_{i}$, divided into a grid of $6400\times 1600$ cells. The ion-to-electron mass ratio is $m_{i}/m_{e}=100$, and the reduced speed of light is $c/V_{A0}=20$. The time step is $\Delta t=0.5\omega_{pe}^{-1}=0.0025\omega_{ci}^{-1}$. The inflow density $n_{c}$ is represented by $\sim 30$ particles per cell, and the total number of particles is $\sim 4.8\times 10^{8}$. The simulation domain has periodic boundary conditions at $x/d_{i}=0$ and $x/d_{i}=L_{x}$, and reflective boundary conditions at $z/d_{i}=-L_{z}/2$ and $z/d_{i}=L_{z}/2$.
@@ IV.2.2 Initial equilibrium
[…]
@@ IV.2.3 Results
Figure 5: Anisotropy $T_{i\bot}/T_{i\parallel}$ (a) and adiabaticity parameter $\kappa$ (b) of initially cold ions in the PIC-HC simulation. The black contour indicates $\kappa=1$. The black, thick-dashed lines indicate the separatrices. The green contour indicates the region of interest for Fig. 4(d).
We present the anisotropy $T_{i\bot}/T_{i\parallel}$ of the initially cold ions at time $t=120~\omega_{ci}^{-1}$ in Fig. 5. At the time and place of investigation, the effect of the periodicity of the simulation is negligible, and the reconnection exhaust is completely dominated by the initially cold ions [55]. Similar to that seen in Fig. 3(j), the temperature anisotropy is mainly parallel outside the current sheet, except in the inflow near the X-line, near the dipolarization front, and at ion mirror points. We compute the adiabaticity parameter $\kappa$ [Fig. 5(b)], and overlay the contour of $\kappa=1$ in Fig. 5(a). We find that, similar to that in the PIC-LP run [Fig. 3(i)], the anisotropy tends to be larger in regions where $\kappa\lesssim 1$ than in the vicinity outside of the current sheet where $\kappa>1$.
We present the joint PDF of $\beta_{i\parallel}$ and $T_{i\bot}/T_{i\parallel}$ in Fig. 4(d). Similar to the PIC-LP run and MMS observations, we restrict to regions of the outflow where $\beta_{i}>0.1$ and $V_{ix}>0$ (green contour in Fig. 5). Compared to the PIC-LP simulation, $T_{i\perp}/T_{i\parallel}$ spans a broader range. This is due to the inclusion of cold ions in the reconnection inflow, which undergo strong (larger than their initial temperature) acceleration in the reconnection outflow, leading to a pronounced parallel temperature anisotropy. In addition, the ion plasma beta remains small $\beta_{i\parallel}\lesssim 10$ in this simulation, allowing a broader range of temperature anisotropies than in higher-$\beta_{i\parallel}$ configurations, such as those inferred from ARTEMIS observations. Nevertheless, similar to MMS observations, the instability thresholds constrain the distribution relatively well for $T_{i\perp}/T_{i\parallel}<1$. In contrast, the perpendicular anisotropy slightly exceeds the marginal stability condition $\kappa=1$. However, similar to MMS observations, this may be attributed to overestimates of $\eta_{e}$ and $\eta_{s}$, or weak scattering, i.e., slow growth of the instability.
@@ V Discussion
We derived two limits for the ion anisotropy in a magnetotail-like quasi-1D current sheet with a finite normal magnetic field, incorporating three ion populations: an adiabatic cold population of two counter-streaming beams, an isotropic hot population, and a quasi-adiabatic Speiser population. Here, we focused on current sheets in the reconnection outflow. Nevertheless, the assumptions of the model with a thin current sheet geometry, finite $B_{z}$, and the superposition of three ion populations are also commonly observed in quiet thin current sheets in the magnetotail [8], suggesting that the derived anisotropy limits are applicable in these configurations as well. The resulting limits define a maximum ion temperature anisotropy as a function of $\beta_{i\parallel}$, similar to the empirical thresholds associated with mirror, firehose, and proton cyclotron instabilities [78]. Unlike these instability criteria, however, the present limits arise from nonlinear particle dynamics within the current sheet, and may constrain the accessible anisotropy prior to the development of kinetic instabilities [65]. Here, we discuss the physical interpretation of these limits and their implications for current sheet stability.
@@
The first limit [Eq. 7], which concerns the parallel anisotropy, resembles the firehose condition, in which the plasma pressure force balances the magnetic tension. In the particular case where $\eta_{c}=1$, i.e., $T_{c\bot}=T_{i\parallel}$ so that the ion VDF consists of a single population with anisotropy $T_{i\bot}/T_{i\parallel}$, the threshold reduces to the well-known fluid firehose condition [25]. In contrast, in the limit $\eta_{c}^{2}\beta_{i\parallel}\rightarrow 0$, the threshold becomes $\mathcal{R}_{c}(\eta_{c}\rightarrow 0,\beta_{i\parallel})\rightarrow\eta_{c}^{2}\beta_{i\parallel}$. This regime corresponds to strongly magnetized cold ions, for which $T_{c\bot}/B^{2}\ll 1$, implying that larger magnetic field curvature is required to scatter the cold ions. Consequently, cold ions can maintain force balance even in highly stretched magnetic field configurations.
@@
The second limit [Eq. 11], which concerns the perpendicular anisotropy, defines the threshold above which the hot isotropic ion population is scattered into quasi-adiabatic transient Speiser orbits. Since the adiabaticity parameter for one particle inversely scales with the particle’s perpendicular velocity $\kappa\propto 1/\sqrt{v_{\bot}}$, for a given increase of the magnetic curvature due to e.g., current sheet thinning/stretching, the supra-thermal ions with the corresponding perpendicular velocity, initially in the hot isotropic background, will be scattered into Speiser orbits. As a result, the number density $n_{h}$ and temperature $T_{h}$ of the hot isotropic background will decrease, and concurrently, the number density $n_{s}$ of the quasi-adiabatic Speiser population will increase, thereby increasing their current density. Eventually, the resulting ion VDF will be composed of a core plus a dense drifting beam Speiser population. If the drift velocity is sufficiently large compared with the thermal speed and the drifting beam is sufficiently dense with respect to the core, this will lead to the growth of a drift-kink type instability [45, 28, 80].
@@
The drift-kink instability may act back on the ion VDFs, scattering quasi-adiabatic Speiser ions into the hot isotropic population. Recent in situ observations suggest that the growth of the drift-kink instability may ultimately lead to a thickening of the current sheet [63]. During the early stage of the drift-kink instability, the enhancement of the quasi-adiabatic Speiser ion current causes the current sheet to thin, generating strong density gradients. These density gradients result in the rapid growth of the lower-hybrid drift instability (LHDI) [29, 30, 27], which is localized at the current sheet edges due to the strong stabilization effect of finite $\beta$ [31]. As demonstrated by Ref. 29 and 27, the nonlinear evolution of the LHDI induces resonant scattering of crossing Speiser ions into the non-crossing region of phase space. This process leads to a net negative charge within the current sheet and the formation of an electrostatic potential across it, which in turn results in a thickening of the current sheet [70]. As the radius of curvature increases due to the thickening of the current sheet, Speiser ions are subsequently scattered back into the hot, isotropic ion population, thereby relaxing the ion VDFs toward equilibrium.
@@
We show that, to maintain the stability of the current sheet, the ion anisotropy is constrained by two limits set by curvature scattering. Nevertheless, as recent theoretical developments and in situ observations suggest, plasma anisotropy and agyrotropy play a crucial role in the force balance of magnetotail-like current sheets [12, 3, 34]. Since electrons are generally isotropic, their pressure force accounts for only up to 10%-30% of the magnetic tension [7]. Recent numerical simulations have further revealed that even a small population of Speiser ions can contribute significantly to maintaining force balance through their anisotropic and agyrotropic pressure contribution [6]. Therefore, our results indicate that Speiser ions play a crucial role in maintaining the force balance of the current sheet and consistently determine the magnetic field curvature, i.e., the degree of stretching of the current sheet, necessary to ensure its stability.